Calculate the pH of the solution that was formed after mixing of 150 mL of 0.200 M H3PO4 with 150 mL of 0.266 M NaOH. Remember to use the Henderson-Hassselbach equation and appropriate pKa value choosing one from pK1=2.12, pK2=7.21, pK3=12.32
1
Expert's answer
2020-05-27T13:35:07-0400

When 150 mL of 0.200 M H3PO4 reacts with 150 mL of 0.266 M NaOH, H3PO4 is converted into NaH2PO4 (0.200-0.066 = 0.134 M) and into Na2HPO4(0.066 M). Therefore, a buffer solution is formed, the buffering action of which is based on the following reaction of dissociation of H2PO4- (conjugate acid) into HPO42- (conjugate base) and H+:

H2PO4- "\rightarrow" HPO42- + H+.

The appropriate pKa value is then pKa2=7.21. According to the Henderson-Hasselbach equation, the pH of a buffer solution is:

"pH = pK_a + \text{log}\frac{[HPO_4^{2-}]}{[H_2PO_4^-]} = 7.21 + \text{log}\frac{0.066}{0.134} = 6.90"

Answer: the pH of the solution is 6.90.